Community gym charges a $50 membership fee and a $55 monthly fee. without gym charges $200 membership fee and a $45 money fee. After how many months will the total amount of money paid to both gyms be the same? What will the amount be?
Found 2 solutions by fractalier, addingup:
Answer by fractalier(6550)

Call x the number of months...then we have
50 + 55x = 200 + 45x
10x = 150
x = 15 months
------------
50 + 55(15) = $875

Answer by addingup(3677)

50+55x = 200+45x
10x = 150
x = 15 months